OpenStudy (anonymous):

List the sides of ΔABC in order from shortest to longest, when m∠A = 10x – 7, m∠B = 7x – 10, and m∠C = 47 – 2x.

jimthompson5910 (jim_thompson5910):

m<A + m<B + m<C = 180 (10x-7) + (7x-10) + (47-2x) = 180 10x-7 + 7x-10 + 47-2x = 180 (10x+7x-2x) + (-7-10+47) = 180 15x + 30 = 180 15x + 30-30 = 180-30 15x = 150 15x/15 = 150/15 x = 10 So m<A = 10x-7 = 10(10)-7 = 93 degree m<B = 7x – 10 = 7(10)-10 = 60 degrees m<C = 47 – 2x = 47-2(10) = 27 degrees Therefore, the side BC is the longest (since it's opposite angle A, which is the largest angle), followed by side AC (it's opposite angle B) and finally side AB is the shortest A pic might help see this.

OpenStudy (anonymous):

this helps alot man! thanks